# Debt, Growth, and the Carbon Lock-In

### v1

@@ Abstract
Despite decades of climate policy and technological progress, global CO2 emissions continue to rise while natural carbon sinks weaken. Here we provide a structural explanation for this persistent overshoot by identifying a link between financial leverage and emissions. We develop a macro-financial model that connects credit dynamics, economic growth, and cumulative carbon emissions. The model extends Kelly’s information-theoretic growth framework to incorporate debt, bankruptcy risk, and emissions trajectories, showing that borrowing amplifies GDP growth and cumulative emissions, thereby locking economies into high-carbon pathways even as energy efficiency improves.
This approach complements traditional Integrated Assessment Models (IAMs) by explicitly linking financial dynamics to cumulative emissions, enabling assessment of financial policy alignment with net-zero pathways. We show that debt increases short-term production gains but at the cost of higher bankruptcy risk and higher cumulative emissions. The model reveals a double constraint: financially, debt requires growth to be repaid; ecologically, growth requires energy, which in turn entails emissions. When intrinsic growth falls below the interest rate, solvency collapses, identifying an optimal leverage frontier beyond which additional debt no longer fuels real wealth and only increases climatic and financial risk.
We calibrate the model using multi-decade macroeconomic and emissions data for the United States, China, France, and Denmark. Across all cases, debt levels, cumulative GDP, and cumulative emissions remain tightly correlated, challenging the feasibility of growth–emissions decoupling and indicating that net-zero targets cannot be achieved without aligning credit allocation with decarbonisation objectives.
In our current financial system, credit expansion amplifies GDP growth and associated emissions, thereby locking economies into higher cumulative carbon trajectories, even as energy efficiency improves through innovation.
@@ Introduction
Debt is a social construct [1, 2], which may appear, for this reason, disconnected from the physical reality. However, debt has real-world implications because borrowed money facilitates the growth of businesses by enabling the consumption of energy and resources, which may be available in finite amounts. Alternatively, it can also cause financial instabilities due to leverage cycles  [3]. Thus, debt brings consumption and production forward by enabling energy investment in resource-intensive sectors, including marginal resource extraction and high-cost infrastructure [2].
In fact, debt creates new feedback loops such as the one represented in Fig. 1, in which credit expansion stimulates growth, which in turn requires additional energy (often directed towards carbon-intensive sectors) and results in carbon emissions, despite ongoing efforts towards decarbonization. In the literature, such a feedback loop is related to the so-called ”macro-economic rebound effect” when investment in clean energy drives productivity increases, which drives growth, thereby also increasing emissions [4].
@@
Motivated by this observation, we study in this paper the link between finance and climate change, focusing particularly on the role of debt. Another motivation for this study is that standard Integrated Assessment Models (IAMs), which typically represent emissions as a function of exogenous GDP growth  [14, 15], do not model the financial system nor investors’ decisions. Thus, the feedback loop represented in Fig. 1 that couples the financial system and mitigation pathways is typically not taken into account by IAMs. Several authors have identified this shortcoming and have criticized IAMs for their lack of financial system [16, 17].
@@
To model the financial system, we rely on Kelly’s gambling model [18], because this model offers a general theory of money, which is not based on notions of economic equilibrium. In 1956, Kelly found a deep connection between information theory and gambling in horse races. This work then became highly influential in investment science [19, 20] and in evolutionary biology [21]. Kelly’s optimal strategy defines an optimal capital allocation by maximizing the expected logarithmic growth rate. The model demonstrates that diversification of assets enables a mitigation of risk from environmental fluctuations, a goal that is also central in biology and ecology and is related to concepts of bet-hedging [11, 21, 22]. In this analogy, the growth of the gambler’s capital is equivalent to the growth of the population, and the risk of bankruptcy is comparable to the extinction of the population. Despite diversification, Kelly’s optimal strategy is known to be risky and this observation hints at a general trade-off between the risk of bankruptcy and the average long-term growth rate of the capital [23, 24].
@@
The original Kelly’s model, however, does not include the possibility of bankruptcy nor the possibility of borrowing(debt). To model these effects, we extend Kelly’s framework to include default dynamics under compounding interest obligations. Once the borrower has taken out a loan, we assume that the debt accumulates with an interest rate at each time step. When the better pays back the debt, the capital is reset at a lower value. The evolution of the capital can thus be compared to a biased random walk with resetting [25, 26, 27], or to a coupled dynamics of money and anti-money variables [28].
@@
In this study, we will use this model as a basis to describe the critical role that credit expansion plays in driving economic growth and cumulative CO2 emissions. Besides quantifying this effect, our stochastic framework also establishes an optimal leverage frontier: the probability of long-term solvency approaches zero when the intrinsic growth rate (growth that originates within an economy, fuelled by factors such as productivity improvements) falls below the effective interest rate, suggesting that structural financial instability is the ultimate effect of excessive borrowing. In this work, we show that credit is the primary driver of several major economies and that this comes at the cost of increasing global dissipation in the form of emissions  [29, 17]. We used country-level economic data to predict the cumulative emissions per capita for these countries. Cumulative CO2 emissions, rather than CO2 emissions, are a key quantity that captures the effect of economic activities in increasing global warming. We demonstrate that to fuel economic growth, credit is required, which has a cost in terms of CO2 emissions.
@@
We present now our extension of Kelly’s model, and we explain in SI how known models of public debt  [33] can be mapped on this framework in sections 1.D and 1.E. Let us assume that at each time step $\tau$ (we use discrete time), the gambler can borrow an amount $B_{\tau}$ with interest rate $\rho_{\tau}$ (real interest rate accounting for inflation [34]). The interest rate $\rho_{\tau}$ is the factor by which the amount borrowed is multiplied at each time step, in other words, the debt due to the amount borrowed at time $\tau^{\prime}$, is multiplied by $\rho_{\tau^{\prime}}^{\tau-\tau^{\prime}}$ if it has not been paid back at time $\tau$. In general, the interest rate is larger than 11 ($\rho>1$), so that the debt keeps increasing, and the borrower ends up paying more than the amount borrowed. Then the gross capital $C_{\tau+1}$, and debt $D_{\tau+1}$ after one step are (until bankruptcy occurs if $C_{\tau+1}\leq 0$):
$C_{0}$ $=B_{0}+C_{\text{ini}}=LC_{\text{ini}},$
$C_{\tau+1}$ $=\gamma_{\tau}C_{\tau}+B_{\tau+1}-\left(\sum_{\tau^{\prime}=0}^{\tau}\delta_{\tau,t_{p}(\tau^{\prime})}\right)D_{\tau+1},$
$D_{0}$ $=B_{0}=(L-1)C_{\text{ini}},$
$D_{\tau}$ $=\sum_{\tau^{\prime}=0}^{\tau-1}\Theta(t_{p}(\tau^{\prime})-\tau)\rho_{\tau^{\prime}}^{\tau-\tau^{\prime}}B_{\tau^{\prime}},$
where $\gamma_{\tau}$ (also accounting for inflation [34]) is the stochastic instantaneous growth rate and depends on time (in a classical Kelly’s model for horse races this factor would equal the product of the bets and odds on the winning horse [23, 24]). The first term in the first equation represents the standard evolution in Kelly’s model, namely the contribution of gambling at time $\tau$: if $\gamma_{\tau}>1$, the gross capital increases provided that there is no borrowing or payback at this time. The second term represents the amount borrowed at time $\tau$; it can be 0 if the borrower does not borrow at this step, or a positive value. The third term represents the possible payback of the debt (debt clearing) at time $\tau+1$, and will either be equal to 0 (no payback), or to $D_{\tau+1}$. $C_{\text{ini}}$ is the amount invested initially by the gambler, and $B_{0}$ is the initial amount borrowed. As usual, the leverage is the ratio of the asset value to the cash needed to purchase it, namely $L=(B_{0}+C_{\text{ini}})/C_{\text{ini}}$ [3].
In addition, $t_{p}(\tau^{\prime})$ indicates the time when the amount borrowed at $\tau^{\prime}$ is paid back. The Heaviside function $\Theta$ function allows us to know if the payback has occurred at time $\tau$ (indeed $\Theta(t_{p}(\tau^{\prime})-\tau)=0$ if $\tau>t_{p}$, meaning that the debt is erased after it has been payed back). If a payback happens at time $\tau$, it means that $\exists\tau^{\prime},\tau=t_{p}(\tau^{\prime})$, and thus $\left(\sum_{\tau^{\prime}}\delta_{\tau,t_{p}(\tau^{\prime})}\right)=1$, otherwise $\left(\sum_{\tau^{\prime}}\delta_{\tau,t_{p}(\tau^{\prime})}\right)=0$.
Without debt, the gross capital cannot hit 0 in Kelly’s model, and bankruptcy never occurs. In our updated version of this model, the capital can become negative if the debt exceeds the capital at the time the debt is repaid. Therefore solvency means that the debt is always paid back at a time where the gross capital is larger ($\gamma_{\tau}C_{\tau}>D_{\tau+1}$). We call $\gamma_{\tau}C_{\tau}-D_{\tau+1}$ the net capital, whose sign indicates if bankruptcy will happen in case of payback. Empirically, the estimated leverage $L_{t}$ remains fairly stable within each economy, varying by only a few percent across years (as shown in Fig. S1 in the SI). This supports the hypothesis of quasi-stationary financial amplification of growth. In the following sections, we outline two different strategies that represent two distinct real-world debt-financed growth regimes.
Finite leverage — initial investment (strategy A)
@@
We assume that the gambler initially borrows money ($B_{0}>0$) but never borrows afterwards, $\forall\tau\geq 1,B_{\tau}=0$. We refer to this strategy as strategy A, which is akin to an individual taking out a loan to buy a house. In practice, the gambler can borrow $B_{0}$ and invest $C_{\text{ini}}$ from its personal capital, and the leverage is then $L=(B_{0}+C_{\text{ini}})/C_{\text{ini}}$. In this case, for $\tau<t_{p}$, the gross capital evolves as in the standard Kelly’s model with an initial investment of $B_{0}+C_{\text{ini}}$, and the debt evolves with an interest rate $\rho$. At the time of payback, the capital can become negative, which could not happen within the purely multiplicative model. Bankruptcy occurs if $C_{t_{p}}$ is negative, or equivalently if $\prod_{\tau^{\prime}\leq t_{p}}\gamma_{\tau^{\prime}}<1-\frac{1}{L}$, meaning that the larger $L$ is, the more likely the bankruptcy. Under the assumption of no bankruptcy, we can also define the long-term growth rate $W=\mathbb{E}\left({\ln\left(\gamma\right)}\right)$. Note that due to the concavity of the logarithm, we have generally for any probability distribution of $\gamma$,
$\ln(E)\geq W,$ (1)
with $E=\mathbb{E}\left({\gamma}\right)$. However, there is a non-zero probability of hitting bankruptcy when the debt is paid back (at $t_{p}$).
Persistent leverage — borrowing at each time step (strategy B)
@@
Let us instead assume that the gambler borrows money at each time step, with leverage $L_{\tau}$ (now time-dependent), which we refer to as strategy B. This strategy is rather that of a state taking yearly on credit to ensure its budget balance. In this case, for $\tau<t_{p}$, the gross capital evolves as in the standard Kelly’s model, with a modified stochastic growth rate, and the debt still evolves with an interest rate $\rho$. The dynamics are then (except for the time of payback)
Now, in addition to the long-term growth rate $W$ defined as before, one needs to also consider the short-term growth rate $\omega(t_{p})$ (defined in SI 1. B). We can separate the value added from production, and that which is due to borrowing in the evolution equations, and define the stochastic intrinsic growth rate as the growth rate to which the part due to borrowing [33] is subtracted:
@@ Results
Now, we scale up the above financial model with strategy B to describe the wealth and the carbon emissions of a country. Let $C_{\tau}$ represents the economic wealth in the year $\tau$ of that country, and let us assume that an external bank lends money each year to that country. As a result, its debt, encapsulated by the variable $D_{\tau}$ increases. Let us also assume that the leverage is constant within the time frame considered. This assumption is justified by the analysis of economic data shown in Fig. S1.
In practice, the GDP during a given period is often used as a measure of economic wealth and has also been criticized for this reason [35, 36]. It consists of summing the monetary values of all products created; therefore, the added growth product between time $\tau$ and $\tau+1$ is $C_{\tau+1}-C_{\tau}$. This capital will be invested in activities that require some energy and produce carbon emissions at time $\tau$, described by the function $\epsilon(\tau,C_{\tau})$. This should be an increasing function of capital, and thus its average should increase with leverage (i.e., the ratio of total money invested to money borrowed). We ultimately have that $\epsilon$ is an increasing function of the capital $C_{\tau}$, which is observed in practice [37].
A key quantity is the carbon intensity of the economy $I(\tau)$ [38], representing the carbon cost of one unit of $GDP$. We connect the emissions to the carbon intensity of the economy thanks to the Kaya relationship [38, 39, 14, 40]:
$\epsilon(\tau,C_{\tau})=C_{\tau}\times I(\tau).$ (4)
The parameter $I(\tau)$ is typically a decreasing function of $\tau$ [38, 14, 40] because energy mixes tend to incorporate more renewable energies over time. Note that $I(\tau)$ can decrease for two main reasons, as illustrated in Fig.1, either the energy cost of $1USD$ decreases, meaning that money is used in less energy-intensive activities, or the carbon cost of energy decreases, meaning that the energy mix of a country includes ”greener” energies. However, this does not mean that cumulative emissions are reduced, as we will see.
@@
We show in SI 2.C that for any strictly decreasing carbon intensities $I(\tau)$, $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ is also a decreasing function of time. We then study the influence of the decarbonization rate (or efficiency of innovations) in SI 2.C and conclude that if $I(\tau)$ is decreasing linearly, the difference between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ stabilizes to a finite value. In general, the difference goes like $(dI/d\tau)/(e^{r}-1)$ at large times. If it decreases faster than linearly, the difference tends to $\infty$, whereas if it decreases more slowly than linearly, the difference first increases and then tends to 0. In practice, it means that the faster decarbonization occurs, the more early carbon-intensive years will matter in the path-dependent intensity. On the contrary, for slow decarbonization, recent years matter more in the path-dependent intensity, making it closer to the carbon intensity.
@@
The larger the path-dependent intensity, the greater the lock-in effect, because we cannot reset $CO_{2}$ emissions to 0 every year. We show path-dependent intensities, as well as intensities in Fig.2. We observe that the orders of magnitude differ widely among countries. We also observe that even if carbon intensities decrease over time, path-dependent intensities remain practically unchanged over a 20-year period. This means that cumulative emissions are largely proportional to cumulative GDP, despite efforts to reduce the annual carbon intensity.
@@
The difference between annual carbon intensity and path-dependent intensity is detailed in the SI. In Fig.2C, we show the correlation between cumulative emissions per capita (since 1998) and cumulative GDP (since 1998) for each country. What we find is that for a carbon intensity decreasing with time and a growing economy, $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ is larger than $I(\tau)$. In particular, the difference between the two depends on the long-term growth rate $W$ and the associated standard deviation $\sigma_{X}$. We can notice that even under steady decarbonization, the ratio of cumulative emissions to cumulative GDP will remain higher than the annual carbon intensity. We also show in SI that larger fluctuations in the economy (larger $\sigma_{X}$) lead in average to a larger decoupling between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ because of phases of economic upturns and high carbon intensities followed by phases of recessions and low carbon intensities. In this case, reducing the annual carbon intensity has little effect on the path-dependent carbon intensity, yielding a larger degree of decoupling.
@@
The trends for $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ are the same across all countries, with varying carbon intensities. For small GDP per capita, the relationship is almost linear and tends to a power law with an exponent slightly smaller than one at larger GDP per capita, as shown in Fig.2C. The lower this exponent, the more significant the efforts to decarbonize the economy.
@@
Figure 2: A Path-dependent intensities $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ starting from 2000 and yearly carbon intensities $I(\tau)$ (consumption-based) for different countries. Data for path-dependent intensities and yearly carbon intensities are shown in dotted lines, theory for these quantities for which $C_{\tau}$ was computed from our model, and $I_{\tau}$ was obtained from data and is shown as a solid line (average), with the colored area representing standard deviation. Even if yearly carbon intensities tend to decrease (the carbon cost of $1USD$ invested decreases with time), path-dependent intensities decrease more slowly than intensities. Therefore, path-dependent intensities can almost be considered as constant on a yearly timescale. In practice, we use data on carbon intensities and our model, which incorporates country-specific measured values of the intrinsic growth rate, interest rate, and leverage, to estimate $C_{\tau}$. We report the mean average errors (MAE) and coefficients of determination between theory and data (path-dependent intensities) for all countries in the Methods section. B Comparing cumulative $CO_{2}$ emissions (data in dotted lines) for different countries (US, China, Denmark, and France) to the results of our stochastic model (in full lines, with standard deviation represented by the colored area). We report the mean average errors (MAE) and coefficients of determination between the theory and data (cumulative emissions per capita since 1998) for all countries in the Methods section. C Relationship between cumulative emissions per capita and cumulative GDP per capita from the data: we observe a clear increasing trend between both, with a slope that varies by country. Here, too, it is clear that China has the most carbon-intensive economy. Yet the US is responsible for the highest cumulative emissions, owing to its high cumulative GDP per capita. D Correlation between cumulative GDP and cumulative debt (left axis), and correlation between cumulative emissions and cumulative debt (right) axis. Cumulative GDP (circles) is mostly linear as a function of cumulative debt, except for Denmark, where the relationship between cumulative GDP and cumulative debt is slightly faster than linear. Concerning cumulative emissions (pentagons), the trend is initially linearly correlated with cumulative debt, then the correlation becomes slower than linear (in particular for China). Again, this is an effect of the decrease in carbon intensity with time.
@@
While the public deficit can stimulate economic activity within a Keynesian framework [46, 47], it also increases cumulative carbon emissions in proportion to the path-dependent carbon intensity. This highlights that even growth-stimulating fiscal policies contribute to emissions lock-in unless accompanied by structural reductions in carbon intensity, and raises the need to question growth targets themselves.
@@
Study of the solvency probability — the likelihood that debt can be paid back before capital depletion
We can relate the probability of bankruptcy to the probability that a biased random walk is smaller than a threshold at payback time $t_{p}$ [52, 53, 25] (the jumps do not have a zero mean, and the random walk is thus biased). We emphasize that bankruptcy does not correspond to the existence of a time where the debt is larger than the gross capital (or first passage time), but to the fact that at the time of payback $t_{p}$, the debt is larger than the gross capital.
@@
We find a probability of solvency.
$P_{S}(t_{p})=\frac{1}{2}\left(1+\text{erf}\left(\frac{t_{p}m-\ln\left(1-\frac{1}{L}\right)}{\sqrt{2t_{p}}\sigma_{X}}\right)\right),$ (9)
where $m=\mathbb{E}(\ln(\gamma/\rho))=W-\ln(\rho)$. We plot this probability on Fig.2A, in the limit of large leverage (associated with higher risk). Interestingly, the result will strongly depend on the sign of $m$:
  • If $m>0$, $P_{S}(t_{p}\to\infty)\to 1$, faster as $\sigma_{X}$ decreases.
  • If $m<0$, $P_{S}(t_{p}\to\infty)\to 0$, faster as $\sigma_{X}$ decreases. In this case, the best course of action is to pay back as soon as possible.
  • If $m=0$, $P_{S}(t_{p}\to\infty)=\frac{1}{2}$.

### v2

@@ Abstract
Despite decades of climate policy and rapid improvements in energy efficiency, global CO2 emissions continue to rise, suggesting the presence of structural drivers that offset efficiency gains. Here we identify financial leverage as a key mechanism underpinning this persistent overshoot.
We develop a stochastic macro-financial model linking credit dynamics, economic growth, bankruptcy risk, and cumulative carbon emissions. The model shows that debt-financed growth systematically amplifies cumulative emissions, locking economies into high-carbon trajectories even as emissions intensity declines. This arises from a double constraint: debt repayment requires sustained growth, while growth remains energy-dependent and thus generates emissions. When growth becomes increasingly dependent on leverage, financial instability and cumulative emissions rise, while gains in real wealth diminish, revealing a leverage frontier beyond which additional credit primarily generates risk.
Calibrating the model to multi-decade data for the United States, China, France and Denmark, we find a robust coupling between debt accumulation, cumulative GDP and cumulative emissions across distinct economic structures. These results challenge the feasibility of growth–emissions decoupling under prevailing credit-driven growth regimes and indicate that achieving net-zero targets requires aligning credit allocation with decarbonisation objectives.
@@ Introduction
Debt is a social construct [1, 2], and may thus appear disconnected from the physical reality. However, debt has real-world implications because borrowed money facilitates the growth of businesses including military activities by enabling the consumption of energy and resources. At the same time, it can also cause financial instabilities due to leverage cycles  [3, 4, 5, 6, 7]. Thus, debt brings consumption and production forward by enabling energy investment in resource-intensive sectors, including marginal resource extraction and high-cost infrastructure [2].
In fact, debt creates new feedback loops in an economy as illustrated in Fig. 1. In such a loop, credit expansion stimulates growth, thereby increasing energy demand [8, 9] (often directed towards carbon-intensive sectors) and results in carbon emissions, despite ongoing efforts towards decarbonization. In the literature, this feedback loop is known as the ’macroeconomic rebound effect’ [10], when investments in clean energy boost productivity and growth, but also increase overall emissions as a result.
@@
Motivated by this observation, we study in this paper the link between finance and climate change, focusing particularly on the role of debt. Another motivation for this study is that standard Integrated Assessment Models (IAMs), which typically represent emissions as a function of exogenous GDP growth  [18, 19], do not model the financial system nor investors’ decisions. Thus, the feedback loop represented in Fig. 1 that couples the financial system and mitigation pathways is typically not taken into account by IAMs. Several authors have identified this shortcoming and have criticized IAMs for their lack of a financial system [7, 20].
@@
To capture these feedback effects, we build a stochastic credit-risk model by combining deterministic macroeconomic models [21, 22, 4, 6], with ingredients from Kelly’s gambling model [23]. In 1956, Kelly found a deep connection between information theory and gambling in horse races. This work then became highly influential in investment science [24, 25] and in evolutionary biology [26]. The model demonstrates that diversification of assets enables a mitigation of risk from environmental fluctuations, a goal that is also central in biology and ecology and is known under the name of bet-hedging [15, 26, 27]. In this analogy, the growth of the gambler’s capital is equivalent to the growth of the population, and the risk of bankruptcy is comparable to the extinction of the population. Kelly’s optimal strategy is known to be risky, and this observation hints at a general trade-off, the risk-return trade-off, between the risk of bankruptcy and the average long-term growth rate of the capital [28, 29].
@@
The original Kelly’s model, however, does not include the possibility of bankruptcy nor the possibility of borrowing (debt). To address this issue, we build a stochastic credit-risk model, that merges the stochastic nature of growth in Kelly’s model with deterministic credit-risk models [4, 6]. In this way, once the borrower has taken out a loan, the debt accumulates with an interest rate at each time step, while the capital evolves stochastically. When the better pays back the debt, the capital is reset at a lower value. The evolution of the capital can thus be compared to a biased random walk with resetting [30, 31, 32], or to a coupled dynamics of money and anti-money variables [33].
@@
In the following, we use this stochastic credit-risk model to investigate the role of credit expansion in driving economic growth and cumulative CO2 emissions. Besides quantifying this causal relationship, our stochastic framework also establishes an optimal leverage frontier: the probability of long-term solvency approaches zero when the intrinsic growth rate (growth that originates within an economy, fueled by factors such as productivity improvements) falls below the effective interest rate, suggesting that structural financial instability is the ultimate effect of excessive borrowing. We also show that credit is the primary driver of several major economies and that this comes at the cost of increasing global dissipation in the form of emissions  [34, 20]. We used country-level economic data to predict the cumulative emissions per capita for these countries. Cumulative CO2 emissions, rather than CO2 emissions, are a key quantity that captures the effect of economic activities in increasing global warming. We demonstrate that to fuel economic growth, credit is required, which has a cost in terms of CO2 emissions.
@@
We consider an investor who can borrow money at each time step $\tau$ with the interest rate $\rho_{\tau}-1$ (real interest rate accounting for inflation [38]), thus increasing both its capital $C_{\tau}$ and its debt $D_{\tau}$. The amount of borrowed money is quantified by the leverage $L_{\tau}$, which is defined as the ratio of the asset value to the cash needed to purchase it [5]. Since we consider an isolated investor, borrowing to an external actor, the leverage can be larger than $1$, whereas in a closed economic system, the total leverage should be $1$. The evolution of the capital is stochastic as in Kelly’s model [24, 25], and is quantified by the intrinsic growth rate (or capital return) $\gamma_{\tau}-1$. The model does not explicitly describe variables such as labor, employment rate or wages, all of which can exhibit complex dynamics [3, 4, 7]. Instead, it focuses on the coupling between the capital and the debt. At each time step, the investor can pay back their debt, and bankruptcy occurs if the debt exceeds the capital when repayment is due. Therefore solvency means that the debt can always be paid back because ($\gamma_{\tau}C_{\tau}>D_{\tau+1}$). We call $\gamma_{\tau}C_{\tau}-D_{\tau+1}$ the net capital, whose sign indicates if bankruptcy will happen in case of payback. In the following sections, we outline two different strategies that represent distinct real-world debt-financed growth regimes. The mathematical model is further detailed in Methods and in sections 1.4 and 1.5 of SI where we explain how our framework relate to known macro-economical models [3, 6, 4, 21, 22].
Finite leverage — one time borrowing (strategy A)
@@
First, we assume that the investor initially borrows money but never borrows afterwards. We refer to this strategy as strategy A, which is akin to an individual taking out a loan to buy a house. In this case, the gross capital evolves as in the classical Kelly’s model, and the debt evolves with a constant interest rate $\rho-1$. At the time of payback, the capital can become negative, and in that case, bankruptcy occurs. Under the assumption of no bankruptcy, we can define the long-term growth rate as the growth rate of the capital when $\tau\to\infty$. However, there is a non-zero probability of hitting bankruptcy when the debt is paid back, and the larger the leverage, the more likely the bankruptcy in this model.
Persistent leverage — periodic borrowing (strategy B)
@@
Let us instead assume that the investor borrows money at each time step, with leverage $L_{\tau}$ (now time-dependent), which we refer to as strategy B. This strategy is rather that of a state or a company taking on credit yearly to ensure its budget balance. In this case, before payback, the gross capital evolves as in the standard Kelly’s model, with a modified stochastic growth rate, and the debt still evolves with an interest rate $\rho-1$. We can separate the value added from production, and that which is due to borrowing in the dynamics, and define the stochastic intrinsic growth rate as the growth rate to which the contribution of borrowing [21] is subtracted.
@@ Results
Now, we scale up the above model based on strategy B to describe the wealth and the carbon emissions of a given country. Let $C_{\tau}$ represent the economic wealth in the year $\tau$ of that country, and let us assume that an external bank lends money each year to that country. As a result, its debt, encapsulated by the variable $D_{\tau}$, increases. Let us also assume that the leverage is constant within the time frame considered. Empirically, the estimated leverage $L_{\tau}$ remains indeed fairly stable within each economy, varying by only a few percent across years (as shown in Fig. S1 in the Supplementary Information). This supports the hypothesis of a quasi-stationary financial amplification of growth.
In practice, the GDP during a given period is often used as a measure of economic wealth and has also been criticized for this reason [39, 40]. With that in mind, we further assume that this economic wealth represents the capital invested in various industries that consume energy to produce carbon emissions at time $\tau$, described by the function $\epsilon(\tau,C_{\tau})$. This function should be an increasing function of the capital, as observed in practice [41], and its average should increase with leverage (i.e., the ratio of total money invested to money borrowed).
We connect the emissions to the carbon intensity of the economy thanks to the Kaya relationship [42, 43, 18, 44]:
$\epsilon(\tau,C_{\tau})=C_{\tau}\times I(\tau),$ (1)
where $I(\tau)$ is the carbon intensity of the economy, measured by the carbon cost of one unit of $GDP$ [42]. The parameter $I(\tau)$ is typically a decreasing function of $\tau$ [42, 18, 44] because energy mixes tend to incorporate more renewable energies over time. Note that $I(\tau)$ can decrease for two main reasons, as illustrated in Fig. 1, either the energy cost of $1USD$ decreases, meaning that money is used in less energy-intensive activities, or the carbon cost of energy decreases, meaning that the energy mix of a country includes ”greener” energies. However, this does not mean that cumulative emissions are reduced, as we will see.
@@
In SI 2.3, we demonstrate that if carbon intensities $I(\tau)$ strictly decrease, then $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ also decreases over time. We then study the influence of the decarbonization rate (or efficiency of innovations) in SI 2.3 and conclude that if $I(\tau)$ is decreasing linearly, the difference between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ stabilizes to a finite value. If $I(\tau)$ decreases faster than linearly, the difference tends to $\infty$, whereas if it decreases more slowly than linearly, the difference first increases and then tends to 0. In practice, this means that the faster decarbonization occurs, the more early carbon-intensive years will matter in the path-dependent intensity. On the contrary, for slow decarbonization, recent years matter more in the path-dependent intensity, making it closer to the carbon intensity.
@@
A higher path-dependent intensity strengthens the lock-in effect, since $CO_{2}$ emissions cannot be reset to zero annually. We show path-dependent intensities, as well as intensities in Fig. 2. We observe that the orders of magnitude differ widely among countries. We also observe that even if carbon intensities decrease over time, path-dependent intensities remain practically unchanged over a 20-year period. This means that cumulative emissions are largely proportional to cumulative GDP, despite efforts to reduce the annual carbon intensity.
@@
The difference between annual carbon intensity and path-dependent intensity is studied more closely in the Supplementary Information. In Fig. 2C, we show the correlation between cumulative emissions per capita (since 1998) and cumulative GDP (since 1998) for each country. When the carbon intensity decreases with time while the economy grows, $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ is larger than $I(\tau)$. The difference between these two quantities depends on the long-term growth rate $W$ and on the amplitude of the fluctuations in the economy measured by $\sigma_{X}$. We observe that even under steady decarbonization, the ratio of cumulative emissions to cumulative GDP remains higher than the annual carbon intensity. Moreover, larger economic fluctuations - as captured by $\sigma_{X}$ - amplify the difference between $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$. This phenomenon arises because periods of economic expansion, characterized by high carbon intensity, are followed by recessions with lower carbon intensity. In this case, reducing the annual carbon intensity has little effect on the path-dependent carbon intensity, yielding a larger degree of carbon lock-in.
@@
The trends for $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ and $I(\tau)$ are similar across all countries, with varying carbon intensities. For countries with small GDP per capita, the relationship is almost linear but bends as a power law with an exponent slightly smaller than one at larger GDP per capita, as shown in Fig. 2C. The lower this exponent, the more significant the efforts to decarbonize the economy.
@@
Figure 2: A Path-dependent intensities $\mathfrak{I}_{\tau}\left(\{C_{\tau}\}\right)$ starting from year 2000 and yearly carbon intensities $I(\tau)$ (consumption-based) for different countries. Data for path-dependent intensities and yearly carbon intensities are shown as dotted lines. The theoretical prediction obtained by computing $C_{\tau}$ from our model, and combining it with $I_{\tau}$ from data, is shown as the solid line representing the average trend, with the colored area around that line representing the standard deviation. In practice, we used country-specific measured values of the intrinsic growth rate, interest rate, and leverage, to estimate $C_{\tau}$. B Comparison of cumulative $CO_{2}$ emissions (data in dotted lines) for different countries (US, China, Denmark, and France) to the predictions of our stochastic model as solid lines, with standard deviation represented by the colored area. We report the mean average errors (MAE) and coefficients of determination between the theory and data for plot A and B in Supplementary Information. C Relationship between cumulative emissions per capita and cumulative GDP per capita from the data: we observe a clear increasing trend between both, with a slope that varies according to the country. As in Fig A, one observes that that China has the most carbon-intensive economy, while the US has the highest cumulative emissions, owing to its high cumulative GDP per capita. D Correlation between cumulative GDP and cumulative debt (left axis), and correlation between cumulative emissions and cumulative debt (right) axis. Cumulative GDP (circles) is mostly a linear function of cumulative debt, except for Denmark, where it is slightly faster than linear. Concerning cumulative emissions (pentagons), the trend is initially linearly correlated with cumulative debt, then becomes slower than linear (in particular for China). Again, this is an effect of the decrease in carbon intensity with time.
@@
While the public deficit can stimulate economic activity within a Keynesian framework [50, 51] (national income is proportional to government spendings through a Keynesian multiplier [22, 52]), it also increases cumulative carbon emissions in proportion to the path-dependent carbon intensity. This highlights that even growth-stimulating fiscal policies contribute to emissions lock-in unless accompanied by structural reductions in carbon intensity, and raises the need to question growth targets themselves.
@@
Study of the solvency probability
Uncertainties in economy cannot be ignored as they can lead to bankruptcy and a lack of solvency. We can relate the probability of bankruptcy to the probability that a biased random walk is smaller than a threshold at payback time which is fixed $t_{p}$ [57, 58, 30]. We emphasize that bankruptcy does not correspond to the existence of a time where the debt is larger than the gross capital (or first passage time), but to the fact that at the time of payback $t_{p}$, the debt is larger than the gross capital. The solvency probability is a related quantity, which represents the likelihood that debt can be paid back before capital depletion.
@@
We find that the probability of solvency depends mainly on the difference between the long-term growth rate and the logarithm of the interest rate. Interestingly, three regimes are possible depending on the sign of this difference, leading to different optimal strategies. If the long-term growth rate is strictly larger than the logarithm of the interest rate, the probability of solvency goes to $1$ as payback time increases, while the opposite trend is observed (solvency probability going to 0) if the long-term growth is strictly lower than the logarithm of the interest rate. Thus, there is no pressure to payback in the first case, while the payback should occur as early as possible in the second case to avoid bankruptcy. In addition, there is a limiting case, where the solvency probability approaches $1/2$ as the payback time tends to infinity.